OpenStudy (anonymous):

How long does it take $1125 to triple if it is invested at 7% interest, compounded quarterly?

OpenStudy (anonymous):

I don't even know how to set this up.

OpenStudy (anonymous):

Rather, I don't know how to solve it. 3375 = 1125(1+.07/4)^94/t), and solve for t. How do I solve for t?

OpenStudy (anonymous):

\[1125(2)=2250\] So that's the money that will be accumulated after years. \[A=P(1+r/n)^(nt)\] is the formula and a is money accumulated p is principal r is annual rate n is number of time intrest is compounded each year t is number of years. So \[2250=1125(1+0.07/4)^(4t)\] \[2=(1+0.07/4)^(4t)\] \[2=1.0175^(4t)\] Then can you do it yourself from this point on? Try this if it works out!